OpenStudy (anonymous):

18. A rectangle is twice as long as it is wide. If the width of the rectangle measures 2.3 inches, what is the area of the rectangle?

OpenStudy (anonymous):

A=w*h w=2.3 h=2w so 2*2.3 Now just substitute and solve

OpenStudy (anonymous):

so 4.6?

OpenStudy (anonymous):

No solve for A=(2.3)(2*(2.3))

OpenStudy (anonymous):

10.58

OpenStudy (anonymous):

yes